OpenStudy (anonymous):

If the diameter of a circle is the same as the radius of another circle, what is the ration of the circumference of the larger circle to the circumference of the smaller?

OpenStudy (anonymous):

let radius of 1st cicle=2 circum=2*pie*r=4*pie let radius of smll circle=1unit circum=2*pie raio of large/small=4*pie/2*pie=2

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

ration of the circumference of the larger circle to the circumference of the smaller is 2

OpenStudy (anonymous):

thanks man